OpenStudy (anonymous):

Triangle ABC is congruent to triangle DEF. In triangle ABC, side AB measures 13, side BC measures x+13, and side CA measures 16. In triangle DEF, side DE measures 13, side EF measures 3x−20, and side FD measures 16. What equation would help you to solve for the side length of BC and EF? Explain your reasoning using complete sentences

OpenStudy (anonymous):

@DangerousJesse @dan815

OpenStudy (anonymous):

Basically I think all they want to know is what equation would help to solve BC and EF and I have to explain my reasoning

OpenStudy (dangerousjesse):

Triangle ABC is congruent to triangle DEF. In triangle ABC, side AB measures 13, side BC measures x+13, and side CA measures 16. In triangle DEF, side DE measures 13, side EF measures 3x−20, and side FD measures 16. What equation would help you to solve for the side length of BC and EF? Explain your reasoning using complete sentences. side BC measures x+13 side EF measures 3x−20 BC ≅ EF corresponding parts of congruent triangles are congruent (CPCTC) if you need more help than that, go to your teacher. (Yahoo)

OpenStudy (anonymous):



OpenStudy (anonymous):

So what is the equation ? sorry im terrible with math and im doing this online and I don't understand it at all

OpenStudy (anonymous):

Therefore, x+13 = 3x-20 x+13 = 3x-20 -13 -13 x = 3x - 33 -3x -2x = -33 x = 16.5 (16.5)+13 = 3(16.5)-20? 29.5 = 29.5

OpenStudy (anonymous):

is that the equation

OpenStudy (anonymous):

AB = DE BC = EF AC = DF

OpenStudy (anonymous):

so the equation is x+13 = 3x-20

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

@larryboxaplenty how would I explain my reasoning on this, I dont get this.